Q=50-4P 

Q=15+P 

what is the surplus in the market if the government sets a price floor of P=10?  

d) what is the shortage in the market if the government imposes a price ceiling of P=5? 

e) calculate the price elasticities of demand and supply when P=7 

f) calculate the price elasticities of demand and supply when P=10.



1
Expert's answer
2021-10-05T13:53:35-0400

Surplus


"Q" "=" 50"-" 4"("10")"

"Q" "=" 50 "-" 40

"Q" "=" 10


"Q" "=" 15 "+" "("10")"

"Q" "=" 15 "+" 10

"Q" "=" 25


Surplus "=" 25 "-" 10 "=" 15



d")" Shortage


"Q" "=" 50"-" 4"("5")"

"Q" "=" 50 "-" 20

"Q" "=" 30


"Q" "=" 15 "+" "(" 5)

"Q" "=" 15 "+" 5

"Q" "=" 20


Shortage "=" 20 "-" 30 "=" "-" 10


e")" Price elasticities of demand and supply when P=7


"%" Using "(" Arc")" or midpoint method of elasticity


Price elasticity of Supply


"Q_o" "=" 15 "+" "("10)

"Q_o" "=" 15 "+" 10

"Q_o" "=" 25



"Q_1" "=" 15 "+" "("7)

"Q_1" "=" 15 "+" 7

"Q_1" "=" 22


"\%" Change in quantity "=" "\frac{Q_1 -Q_o}{(P_1+p_0)/2}"


Price elasticity of supply "=" "\frac{22-25}{(7+10)/2}"


Price elasticity of supply "=" "\frac{-3}{8.5}" "=" 0.353


Price elasticity or demand



"Q_o" "=" 50"-"4 "("10)

"Q_o" "=" 50 "-" 40

"Q_o" "=" 10



"Q_1" "=" 50 "-"4"("7)

"Q_1" "=" 50 "-" 28

"Q_1" "=" 22


"\%" Change in quantity "=" "\frac{Q_1 -Q_o}{(P_1+p_0)/2}"


Price elasticity of demand "=" "\frac{22-10}{(7+10)/2}"


Price elasticity of demand "=" "\frac{12}{8.5}" "=" 1.411


f) Price elasticity of supply


"Q_o" "=" 15 "+" "("5)

"Q_o" "=" 15 "+" 5

"Q_o" "=" 20



"Q_1" "=" 15 "+" "("10)

"Q_1" "=" 15 "+" 10

"Q_1" "=" 25


"\%" Change in quantity "=" "\frac{Q_1 -Q_o}{(P_1+p_0)/2}"


Price elasticity of supply "=" "\frac{25-20}{(10+5)/2}"


Price elasticity of supply "=" "\frac{5}{7.5}" "=" 0.666


Price elasticity or demand



"Q_o" "=" 50"-"4 "("5)

"Q_o" "=" 50 "-" 20

"Q_o" "=" 30



"Q_1" "=" 50 "-"4"("10)

"Q_1" "=" 50 "-" 40

"Q_1" "=" 10


"\%" Change in quantity "=" "\frac{Q_1 -Q_o}{(P_1+p_0)/2}"


Price elasticity of demand "=" "\frac{10-30}{(10+5)/2}"


Price elasticity of demand "=" "\frac{-20}{7.5}" "=" 2.66